OpenStudy (anonymous):

Solve using elimination. Show ALL work. 2x – 5y = 19 4x + 10y = -22

OpenStudy (anonymous):

@oldrin.bataku help on this one??

OpenStudy (anonymous):

\[ \ \ \ 2x-5y=19\\ \ \ \ 4x+10y=-22\\ \text{Multiply the first by 2.}\\ \ \ \ 4x-10y=38\\ \text{Combine the first and second.}\\ \ \ \ 8x=16\\ \ \ \ x=2\\ \text{Find }y\text{ after solving for }x.\\ \ \ \ 4(2)+10y=-22\\ \ \ \ 10y=-22-8=-30\\ \ \ \ y=-3 \]

OpenStudy (anonymous):

Thanks!!